OpenStudy (anonymous):

Explain how opposites are applied when solving one-step equations. How do you know what to use to solve a one-step equation? Be specific and complete.

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

hello

OpenStudy (misty1212):

what the monkey is a "one step" equation?

OpenStudy (misty1212):

does that mean you can get the answer in "one step"?

OpenStudy (anonymous):

when you only have to do one step to solve it i think

OpenStudy (anonymous):

yeah

OpenStudy (misty1212):

lol some math teacher must have coined that phrase

OpenStudy (anonymous):

I think I know the answer just want to make sure with someone

OpenStudy (misty1212):

ok so it is something like \[x+5=9\]i guess you are supposed to the the "opposite" of \(5\) is \(-5\) so we can solve via \[x+5-5=9-5\\ x=4\]

OpenStudy (anonymous):

yes spot on

OpenStudy (misty1212):

or if you have \[x-4=7\] then the "opposite" of \(-4\) is \(4\) so \[x-4+4=7+4\\ x=11\]

OpenStudy (anonymous):

yeah

OpenStudy (misty1212):

wonder if they taught them this is math education school it is certainly not math that is for sure

OpenStudy (anonymous):

it is

OpenStudy (misty1212):

maybe include \[2x=10\] the "opposite" of multiplying by 2 is dividing by 2, so \[2x=10\\ \frac{2x}{2}=\frac{10}{2}\\ x=5\]

OpenStudy (anonymous):

okay thank you :)

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]